OpenStudy (anonymous):

Choose the equation of the line passing through the point (2, 6) and parallel to y = -3x - 4. y = 3x - 12 y = -3x y = -3x + 12 y = 3x

OpenStudy (anonymous):

Since the line is parallel to y=-3x+4 then the slope must be equal so m=-3. Then just plug into \[y-y_1=m(x-x_1)\] and you'll get \[y-6=-3(x-2)\]Now just simplify and solve for y in order to get your answer.

OpenStudy (johnweldon1993):

slope of a parallel line is the same as the other..so -3 use point slope formula to find this equation y - 6 = -3(x-2) y - 6 = -3x + 6 y = -3x + 12

OpenStudy (johnweldon1993):

And now check.... 6 = -3(2) + 12 6 = -6 + 12 -6 = -6 yes it does so it is correct